Jeremy Pruitt wants to attend an engineering class with his quarterback so he sits in on an EF 157 lecture. He is so impressed with the garrulous use of vectors, vector hats, and coordinate systems that he changes all of the teams pass routes to vectors For their best play, he states that the quarterback throws from the origin while the receiver starts at position x,y- (-15.0,2.0). The receiver runs 30 yds at a heading of 80°, and then 30 yds at a heading of 345° to the point where hell catch the footbal (all distances are in yards, and all angles are measured CCW from the x axis) This problem DOES need to be written out for your portfolio NOTES IMAGES DISCUSS UNITS STATS HELP Part Description Answer Chk History How far does the pass have to go to reach the receiver? (include units with answer) 25 pts, 100% # tries: 0 Show Details Format Check 2% try penalty Hints: 1.2 What direction must the pass be thrown? (include units with answer) 25 pts, 100% # tries: O show Details Format Check 296 try penalty Hints: 0,
Answer #1

Initial location of the receiver P = -15.0 i + 2.0 J

Location of the quarterback after the first movement is

P1 = (-15.0 + 30cos 80)i + (2.0 + 30sin80)j

P1 = -9.8 i + 31.5 j

Location of the quarterback after the second movement is

P2 = (-9.8 + 30cos345 )i + (31.54 + 30sin345)j

P2 = 19.2 i + 23.8 j

P2 is the receiving point

  1. Total distance of the pass is d = √(19.22 + 23.82) = 30.57 = 30.6 yards
  2. direction of the pass θ = tan-1(23.8/19.2)= 51.1o